OpenStudy (crissy15):

Select the expression equal to square root of 52, I'm forgetting simply things, can anyone refresh how this is again?

OpenStudy (crissy15):

4 square root 3 4 square root 13 2 square root 13 13 square root 2

OpenStudy (bibby):

52=4*13 use that and the fact that \(\sqrt{a*b}=\sqrt a * \sqrt b\)

OpenStudy (crissy15):

ah right ok thanks XD